OpenStudy (anonymous):

Eliminate or Addition 3x-y+z=-15 x+2y-z=1 2x+3y-2z=0

OpenStudy (anonymous):

@waterineyes I can't Eliminate x or y.

OpenStudy (anonymous):

I tried but I think I am wrong. I got equation 4 but I cant get equation 5

OpenStudy (anonymous):

Don't go by Substitution.. Do you know Elimination Method??

OpenStudy (anonymous):

yeah but this is kinda hard. I got 4x + y =-14 in my equation 4 but my equation % is 7y=1

OpenStudy (anonymous):

equation 5

OpenStudy (anonymous):

I am not getting what you are saying.. Tell me at very first what you did so I can know what car are you driving to solve this.. Ha haha..

OpenStudy (anonymous):

3x-y+z=-15 + x-2y-z=1 equal equation 4 which is 4x+y=-14

OpenStudy (anonymous):

Oh sorry you are correct..

OpenStudy (anonymous):

x-2y-z=1 multiplied by -2 -2x+2y+2z==0

OpenStudy (anonymous):

Yes you will have equation 4 as : \(4x +y = -14\)

OpenStudy (anonymous):

Is that +2y or -2y in the second equation??

OpenStudy (anonymous):

I think you are right 4x-3y=-14

OpenStudy (anonymous):

Firstly clear my point.. Check the second equation from the text and write it here again..

OpenStudy (anonymous):

Is it 2y or -2y ???

OpenStudy (anonymous):

3x-y+z=-15 + x+2y-z=1 equal equation 4 which is 4x+y=-14

OpenStudy (anonymous):

Yes now you are right.. Now multiply -2 to the second equation and tell me what you got..

OpenStudy (anonymous):

x+2y-z=1(-2) -2x-4y+2z=-2

OpenStudy (anonymous):

Yes now add this with the third equation..

OpenStudy (anonymous):

I think now you got it.. -y = -2 y = 2.. Find x and z then..

OpenStudy (anonymous):

-2x-4y+2z=-2 + 2x+3y-2z=0 equal to -y=-2

OpenStudy (anonymous):

@waterineyes amI right?

OpenStudy (anonymous):

do I need to devide it in -1 to get y=2? @waterineyes

OpenStudy (anonymous):

Yes you are right.. -y = -2 multiply by -1, y = 2.. Now find x and z..

OpenStudy (anonymous):

x=-4 and z=-1?

OpenStudy (anonymous):

Let me check wait..

OpenStudy (anonymous):

Yes dear you are absolutely right...

OpenStudy (anonymous):

Thanks a lot for your help dear

OpenStudy (anonymous):

Welcome nia..